# Analysis of cohort stepped wedge cluster-randomized trials with non-ignorable dropout via joint modeling

### v1

@@ 3 A joint modeling approach to address informative attrition in cohort stepped wedge designs
When dropout from the study is informative, the traditional mixed-effects analysis models introduced in Section 2 can yield biased parameter estimates as it assumes outcomes are MAR, i.e., that missingness is non-informative. To overcome this limitation, we propose jointly modeling the longitudinal and dropout processes via joint longitudinal-survival modeling.
We now introduce additional notation to describe the dropout process. Specifically, let $S_{ik}$ be the full observation time for the kth participant in the ith cluster. This corresponds to the full time under study, e.g., until the end of the jth intervention period, for a patient that completes the study. However, study participants can drop out before the end of the last intervention period for various reasons, including death or migration from the country. Therefore, in practice, we can only observe time $T_{ik}=\min(S_{ik},C_{ik})$, where $C_{ik}$ is the dropout time; the indicator variable $\Delta_{ik}=I(C_{ik}\leq S_{ik})$ represents whether a participant dropped out of the study before the end of follow up ($\Delta_{ik}=1$) or not ($\Delta_{ik}=0$). The observed time to event data will thus be, for the kth participant in the ith cluster, the couple $(T_{ik},\Delta_{ik})$.
@@ 3.1 Joint model formulation
where $Z_{ik}=[\alpha_{i},\phi_{ik}]$ denotes any user-specified combination of the random effects $\alpha$, $\phi$ with associated regression coefficients $\omega$; the dimension of $\omega$ depends on which random effects from the longitudinal submodel are included in $Z_{ik}$. We denote this parametrization of the association structure between the longitudinal and time-to-dropout models as the shared random effects parametrization. Moreover, the survival submodel assumes a proportional hazards model with baseline hazard function $\lambda_{0}(t)$ and treatment effect $\nu$, but other covariates could be incorporated in the dropout model, in principle.
@@
where we allow for cluster- and subject-specific associations with the rate of dropout from the study (with association parameters $\omega_{1},\omega_{2}$). We can simplify this maximal model by including constraints: for instance, we could constrain the effect of cluster- and subject-specific random effects on dropout to be the same (i.e., $\omega_{1}=\omega_{2}$) or that a certain effect is zero (e.g., $\omega_{2}=0$). Any other user-defined constraint is also possible.
@@
Moreover, note that we assume continuous time for the dropout component of the joint model, and assume a parametric shape for the baseline hazard function $\lambda_{0}(t)$, such as exponential or Weibull. For instance, an exponential baseline hazard function has a constant hazard defined by a single parameter $\lambda$, with $\lambda_{0}^{\text{Exp}}(t)=\lambda$, while a Weibull baseline hazard function with scale and shape parameters $\lambda$ and $p$, respectively, is defined as $\lambda_{0}^{\text{Wei}}(t)=\lambda pt^{p-1}$. Other parametric (e.g., Gompertz) or flexible parametric (e.g., using smooth splines to model the baseline hazard) formulations can be accommodated as well within the framework, in principle 29, 30. Semi-parametric and fully non-parametric baseline hazard functions are also possible, in principle, but that can lead to underestimation of the standard errors of parameter estimates requiring the bootstrap to obtain appropriate standard errors 31.
Furthermore, we assume that dropout times are known exactly: in this scenario, we only have to consider right censoring for the time to dropout distribution. In some settings, information that people have dropped out of the study might be only known at the end of each study period; in that scenario, the distribution for time to dropout would be interval censored, and the joint modeling approach would have to be extended to take this into account 32.
@@ 3.2 Likelihood and estimation
where $\theta$ denotes the vector of model parameters, $\mathbf{y}$ the vector of response variables (the longitudinal and time to dropout outcomes), $\mathbf{x}$ the vector of model covariates (such as the fixed treatment and period effects), $f(\cdot|\sigma^{2}_{\alpha})$ the distribution of the cluster-specific random intercept $\alpha$, and $f(\cdot|\sigma^{2}_{\phi})$ the distribution of the subject-specific random intercept $\phi$. The joint distribution of the random effects can be factorized as the product of their univariate distributions thanks to the independence assumption ($\alpha\perp\!\!\!\perp\phi$) that follows from the nested exchangeable model formulation. The integration in Equation (6) does not have closed-form, and needs to be approximated numerically; to this end, we use mean-variance adaptive Gauss-Hermite quadrature as implemented in Stata’s gsem command 33, 34.
A standard assumption in this class of models is that, conditionally on the random effects, the longitudinal measurements are independent of dropout times, i.e., $Y\perp\!\!\!\perp T_{ik}\ |\ \alpha,\phi$ (using the notation of Section 3.1). Therefore, we can factorize the joint distribution of the longitudinal measurements and dropout times as well:
$f(\mathbf{y}|\mathbf{x},\alpha,\phi,\theta)=f(Y|\mathbf{x}^{Y},\alpha,\phi,\theta^{Y})\times f(t|\mathbf{x}^{t},\alpha,\phi,\theta^{t})$ (7)
where $Y$ denotes the longitudinal outcome and $t$ the time to dropout outcome. The vector of fixed effects covariates is partitioned into covariates for the longitudinal submodel ($\mathbf{x}^{Y}$) and for the survival submodel ($\mathbf{x}^{t}$), and analogously the vector of model parameters $\theta=(\theta^{Y},\theta^{t})$. In practice, for our two-level model, the likelihood is computed at the cluster level by multiplying the subject-specific contribution of each subject in each cluster first and then combining all cluster contributions.
The likelihood function does not have a closed-form solution and is thus maximized numerically, e.g., using Stata’s implementation of the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm or its modified Newton-Raphson algorithm 35. We implement the entire model in Stata, including Gaussian quadrature and optimization, within the generalized structural equation modeling framework using the gsem command 34. More details on the estimation of hierarchical joint models can be found in Brilleman et al. 36
@@ 4 Monte Carlo simulation studies
The setup of our simulation studies is described in this Section, following the ADEMP framework introduced by Morris et al. 37
@@ 4.2 Data-generating mechanisms
We assume that the time to dropout submodel follows a Weibull proportional hazards model with shape and scale parameters $p=1.0$ and $\lambda=\exp(-1.5)$. These values are based on the real-data application that is described in Section 5. Then, we assume that the five parameters for the longitudinal period effects are $\beta_{1}=\beta_{2}=\beta_{3}=\beta_{4}=\beta_{5}=30$. For the variance components, we assume a residual error with variance $\sigma^{2}_{\varepsilon}=40$, and random intercepts variances of $\sigma^{2}_{\alpha}=2$ and $\sigma^{2}_{\phi}=55$. These values correspond to a within-individual ICC ($\rho_{a}$) of 0.588 and a between-individuals ICC ($\rho_{d}$) of 0.021 which are within the range of commonly reported correlation parameters in stepped wedge designs 38.
@@
  • The association parameters $\omega_{1}=\omega_{2}\in\{\log(0.5),\log(0.9),\log(1.0),\log(2.0)\}$. These correspond to hazard ratios of $0.5$, $0.9$, $1.0$, and $2.0$, respectively, and we deem scenarios with $\omega_{1}=\omega_{2}\in\{\log(0.5),\log(2.0)\}$ as extremes.
@@
The procedure to simulate data from the joint model was developed by combining previous work by Bender et al., Austin, and Crowther and Lambert 39, 40, 41, and consists of two steps. The first step consists of simulating the longitudinal data from a stepped wedge trial in the absence of dropout; then, we simulate the time to dropout from the study and censor the longitudinal trajectories accordingly. These steps are formalized in Algorithm 1. Simulating dropout times from the survival submodel of Equations (3), (5) follows from Bender et al. 39, and is based on the inversion method. Specifically, let $T^{s}$ be the simulated dropout time; assuming that
@@
with $u\sim\text{Unif}(0,1)$, we can solve for $T^{s}$ and obtain simulated dropout times. However, in our case, this is not straightforward as the cumulative hazard function depends on (1) the time-varying treatment, (2) the time-period-specific effects, and (3) the random effects. We can thus implement the procedure described by Austin 40: assuming that $t_{J}$ is the start of the time interval where a given subject starts to be exposed to treatment, a Weibull baseline hazard with shape and scale parameters $\lambda$ and $p$, respectively, and $b=\lambda\exp(\omega_{1}\alpha_{i}+\omega_{2}\alpha_{ik}){t_{J}}^{p}$, dropout times can be simulated as
@@
with, again, $u\sim\text{Unif}(0,1)$. Similar closed-form expression could be obtained for exponential or Gompertz baseline hazard functions, again, following the work of Austin 40. We provide an R package with our implementation of this algorithm, which is openly available online on GitHub at https://github.com/RedDoorAnalytics/simswjm.
@@ 4.5 Performance measures
The key performance measure of interest was bias, which quantifies whether a method targets the true value of a parameter on average. We report bias on both an absolute and relative scale, the latter only for estimands with a true value that is not zero. Additionally, we report empirical and model-based standard errors to assess the performance of variance estimators and to compare the efficiency of the two methods. Finally, we report convergence rates and coverage probabilities, for completeness. All performance measures are described in more detail by Morris et al. 37
@@ 4.6 Number of repetitions
To determine how many repetitions were needed to estimate the performance measures of interest with sufficient precision, we first ran an arbitrary amount of 30 iterations. We then analyzed these preliminary results and concluded that, with 500 repetitions, we would expect a Monte Carlo error for bias that is no more than 2% of the absolute treatment effect value (either of $\delta$ or $\delta_{0},\delta_{1},\delta_{2},\delta_{3}$), which we consider reasonable. This procedure is described in more detail in the supplementary material.
@@ 4.7 Software
The simulation algorithms were implemented in R, while the analysis code for both models (linear mixed and joint model) was implemented in Stata using the built-in gsem command 42, 34. The simulation study is summarized in R using the rsimsum package 43. All statistical code to replicate our simulations or fit the joint model using Stata is publicly available on GitHub at https://github.com/ellessenne/swjm.
@@ 4.8 Results
We report the results of the simulation study for the constant intervention model in 4.8.1, and the results for the general time on treatment model in Section 4.8.2.
@@ 4.8.1 Results: simulation study A, constant intervention model
First, we report on the convergence rates of the fitting procedures. Overall, the joint model had lower convergence rates than the mixed model, which was expected given the extra computational complexity; nonetheless, convergence rates were good (>97%) in all scenarios with non-extreme dropout mechanisms. These results are described in more detail in Appendix C of the supplementary material available online.
Figure 5 depicts the estimated bias for the treatment effect $\delta$ on the longitudinal outcome over repetitions and across all simulation scenarios and models; results are also tabulated in Appendix C of the supplementary material. Overall, the joint model outperformed the linear mixed model, with negligible bias under every scenario. Conversely, the linear mixed effects model showed bias estimates in the scenarios with $\nu=-0.2$ and informative dropout, irrespective of the true treatment effect value. These biases were, however, of small magnitude, as relative bias did not exceed 1.1% (in absolute terms; Appendix C). In all scenarios with $\nu=0.0$ both models yielded unbiased results, and, as expected, both models performed well in all scenarios with non-informative dropout ($\omega_{1}=\omega_{2}=\log(1.0)$).
@@
Interestingly, coverage probability was close to optimal across all scenarios and for both models, highlighting some issues with the estimation of model-based standard errors for linear mixed models in the settings of informative dropout. Some exceptions could be observed, with slight under-coverage for both models in the scenario with $\omega_{1}=\omega_{2}=\log(1.0),\nu=0.0,\delta=0.0$ and for the linear mixed model in the scenario with $\omega_{1}=\omega_{2}=\log(0.5),\nu=-0.2,\delta=25.0$; we however suspect that these results may be due to the availability of a limited number of clusters, and we suspect that they may improve as the number of clusters increases (the coverage probabilities are still in the 90-95% range). These results are reported in Appendix C of the supplementary material.
For the period effects, the linear mixed model showed significant bias at all possible time points when dropout was informative (Appendix C, supplementary material), up to 20% (on a relative scale); the joint model outperformed the linear mixed model, with slight bias (<5%) only in the extreme scenarios with $\omega_{1}=\omega_{2}=\log(0.5),\log(2.0)$. Accordingly, coverage probability was poor when period coefficients were biased for both the linear mixed model and the joint model, and optimal otherwise. As with the treatment effect $\delta$, when dropout was not informative, both models could estimate unbiased period effects with optimal coverage probabilities.
For the ICC, the joint model outperformed the linear mixed model in all scenarios with informative dropout, showing no bias or small bias (at worst, in some extreme scenarios), as in Figure 6; accordingly, coverage was low (Appendix C, supplementary material). The under-coverage is not unexpected because typically, nominal coverage for such second-order parameters may require a large number of clusters, as demonstrated in previous simulations 44. More results on the estimation of the specific variance components are included, once again, in Appendix C of the supplementary material.

### v3

@@ 4 A joint modeling approach to address informative attrition in cohort stepped wedge designs
When dropout from the study is informative, the traditional mixed-effects analysis models introduced in Section 3 can yield biased parameter estimates as it assumes outcomes are MAR, i.e., that missingness is non-informative. To overcome this limitation, we propose to simultaneously model the longitudinal and dropout processes via shared random-effects longitudinal-survival modeling.
We now introduce additional notation to describe the dropout process. Specifically, let $S_{ik}$ be the full observation time for the kth participant in the ith cluster. This corresponds to the full time under study, e.g., until the end of the final intervention period, for a patient that completes the study. However, study participants can drop out before the end of the last intervention period for various reasons, including death or migration from the country. Therefore, in practice, we can only observe time $T_{ik}=\min(S_{ik},C_{ik})$, where $C_{ik}$ is the dropout time; the indicator variable $\Delta_{ik}=I(C_{ik}\leq S_{ik})$ represents whether a participant dropped out of the study before the end of follow up ($\Delta_{ik}=1$) or not ($\Delta_{ik}=0$). The observed time to event data will thus be, for the kth participant in the ith cluster, the couple $(T_{ik},\Delta_{ik})$.
@@ 4.1 Joint model formulation
where $Z_{ik}=[\alpha_{i},\phi_{ik}]$ denotes any user-specified combination of the random effects $\alpha$, $\phi$ with associated regression coefficients $\omega$; the dimension of $\omega$ depends on which random effects from the longitudinal submodel are included in $Z_{ik}$. Note that the association parameter(s) $\omega$ are sometimes referred to as factor loadings. 36 We denote this parametrization of the association structure between the longitudinal and time-to-dropout models as the shared random-effects parametrization. Moreover, the survival submodel assumes a proportional hazards model with baseline hazard function $\lambda_{0}(t)$ and treatment effect $\nu$, but other covariates could be incorporated in the dropout model, in principle.
@@
where we allow for cluster- and subject-specific associations with the rate of dropout from the study (with association parameters $\omega_{1},\omega_{2}$). We can simplify this maximal model by including constraints: for instance, we could constrain the effect of cluster- and subject-specific random effects on dropout to be the same (i.e., $\omega_{1}=\omega_{2}$) or that a certain effect is zero (e.g., $\omega_{2}=0$).
@@
Moreover, note that we assume continuous time for the dropout component of the joint model, and assume a parametric shape for the baseline hazard function $\lambda_{0}(t)$, such as exponential or Weibull. For instance, an exponential baseline hazard function has a constant hazard defined by a single parameter $\lambda$, with $\lambda_{0}^{\text{Exp}}(t)=\lambda$, while a Weibull baseline hazard function with scale and shape parameters $\lambda$ and $p$, respectively, is defined as $\lambda_{0}^{\text{Wei}}(t)=\lambda pt^{p-1}$. Other parametric (e.g., Gompertz) or flexible parametric (e.g., using smooth splines to model the baseline hazard) formulations can be accommodated as well within the framework, in principle. 37, 38 Semi-parametric and fully non-parametric baseline hazard functions are also possible, in principle, but that can lead to underestimation of the standard errors of parameter estimates requiring the bootstrap to obtain appropriate standard errors. 39 The continuous-time modeling strategy also differentiates this work from that of Wang and Chinchilli, who considered a survival submodel based on a discrete-time formulation. 23
Furthermore, we assume that dropout times are known exactly: in this scenario, we only have to consider right censoring for the time to dropout distribution. In some settings, information that people have dropped out of the study might be only known at the end of each study period; in that scenario, the distribution for time to dropout would be interval censored, and the joint modeling approach would have to be extended to take this into account. 40
@@ 4.2 Likelihood and estimation
where $\theta$ denotes the vector of model parameters, $\mathbf{y}$ the vector of response variables (the longitudinal and time to dropout outcomes), $\mathbf{x}$ the vector of model covariates (such as the fixed treatment and period effects), $f(\cdot|\sigma^{2}_{\alpha})$ the distribution of the cluster-specific random intercept $\alpha$, and $f(\cdot|\sigma^{2}_{\phi})$ the distribution of the subject-specific random intercept $\phi$. The joint distribution of the random effects can be factorized as the product of their univariate distributions thanks to the independence assumption ($\alpha\perp\!\!\!\perp\phi$) that follows from the nested exchangeable model formulation. The integration in Equation (6) does not have closed-form, and needs to be approximated; to this end, we use mean-variance adaptive Gauss-Hermite quadrature as implemented in Stata’s gsem command. 41, 42
A standard assumption in this class of models is that, conditionally on the random effects, the longitudinal measurements are independent of dropout times, i.e., $Y_{ijk}\perp\!\!\!\perp T_{ik}\ |\ \alpha,\phi$ (using the notation of Section 4.1). Therefore, we can factorize the joint distribution of the longitudinal measurements and dropout times as well:
$f(\mathbf{y}|\mathbf{x},\alpha,\phi,\theta)=f(\mathbf{Y}|\mathbf{x}^{Y},\alpha,\phi,\theta^{Y})\times f(\mathbf{t}|\mathbf{x}^{t},\alpha,\phi,\theta^{t})$ (7)
where $\mathbf{Y}$ denotes the longitudinal outcome and $\mathbf{t}$ the time to dropout outcome. The vector of fixed effects covariates is partitioned into covariates for the longitudinal submodel ($\mathbf{x}^{Y}$) and for the survival submodel ($\mathbf{x}^{t}$), and analogously the vector of model parameters $\theta=(\theta^{Y},\theta^{t})$. In practice, for our two-level model, the likelihood is computed at the cluster level by multiplying the subject-specific contribution of each subject in each cluster first and then combining all cluster contributions.
The likelihood function does not have a closed-form solution and is thus maximized numerically, e.g., using Stata’s implementation of the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm or its modified Newton-Raphson algorithm. 43 We implement the entire model in Stata, including Gaussian quadrature and optimization, within the generalized structural equation modeling framework using the gsem command. 42 More details on the estimation of hierarchical joint models can be found in Brilleman et al. 44
@@ 5 Monte Carlo simulation studies
The setup of our simulation studies is described in this Section, following the ADEMP framework introduced by Morris et al. 45
@@ 5.2 Data-generating mechanisms
We assume that the time to dropout submodel follows a Weibull proportional hazards model with shape and scale parameters $p=1.0$ and $\lambda=\exp(-1.5)$. These values are based on the real-data application that is described in Section 6. Then, we assume that the five parameters for the longitudinal period effects are $\beta_{1}=\beta_{2}=\beta_{3}=\beta_{4}=\beta_{5}=30$. For the variance components, we assume a residual error with variance $\sigma^{2}_{\varepsilon}=40$, and random intercepts variances of $\sigma^{2}_{\alpha}=2$ and $\sigma^{2}_{\phi}=55$. These values correspond to a within-individual ICC ($\rho_{a}$) of 0.588 and a between-individuals ICC ($\rho_{d}$) of 0.021 which are within the range of commonly reported correlation parameters in stepped wedge designs. 46
@@
  • The association parameters $\omega_{1}=\omega_{2}\in\{\log(0.5),\log(0.9),\log(1.0),\log(2.0)\}$. These correspond to hazard ratios of $0.5$, $0.9$, $1.0$, and $2.0$, respectively, and we deem scenarios with $\omega_{1}=\omega_{2}\in\{\log(0.5),\log(2.0)\}$ as extremes to demonstrate the performance of different methods in challenging scenarios.
@@
The procedure to simulate data from the joint model was developed by combining previous work by Bender et al., Austin, and Crowther and Lambert, and consists of two steps. 47, 48, 49 The first step consists of simulating the longitudinal data from a stepped wedge trial in the absence of dropout; then, we simulate the time to dropout from the study and censor the longitudinal trajectories accordingly. These steps are formalized in Algorithm 1. Simulating dropout times from the survival submodel of Equations (3), (5) follows from Bender et al., and is based on the inversion method. 47 Specifically, let $T^{s}$ be the simulated dropout time; assuming that
@@
with $u\sim\text{Unif}(0,1)$, we can solve for $T^{s}$ and obtain simulated dropout times. However, in our case, this is not straightforward as the cumulative hazard function depends on (1) the time-varying treatment, (2) the time-period-specific effects, and (3) the random effects. We can thus implement the procedure described by Austin: 48 assuming that $t_{J}$ is the start of the time interval where a given subject starts to be exposed to treatment, a Weibull baseline hazard with shape and scale parameters $\lambda$ and $p$, respectively, and $b=\lambda\exp(\omega_{1}\alpha_{i}+\omega_{2}\phi_{ik}){t_{J}}^{p}$, dropout times can be simulated as
@@
with, again, $u\sim\text{Unif}(0,1)$. Similar closed-form expression could be obtained for exponential or Gompertz baseline hazard functions, again, following the work of Austin. 48 We provide an R package with our implementation of this algorithm, which is openly available online on GitHub at https://github.com/RedDoorAnalytics/simswjm.
@@
Finally, we extend one of the simulation scenarios described above to test a few more settings in a non-factorial fashion (given the extensive number of simulation scenarios already implemented). Specifically, we start with a scenario with informative dropout ($\nu=-0.2,\omega_{1}=\omega_{2}=\log(0.9)$) and with a non-zero treatment effect ($\delta=5.0$). Then, we update it by (1) reducing the number of clusters randomized to each intervention sequence to 3 (instead of 8), (2) reducing to 3 clusters per intervention sequence but doubling the number of participants per cluster to 100, (3) reducing the ICCs values by halving the variance of the cluster- and subject-specific random intercepts ($\sigma^{2}_{\alpha}=1,\sigma^{2}_{\phi}=27.5,\rho_{\alpha}=0.416,\rho_{d}=0.015$), and (4) increasing the ICCs values by doubling the random intercept variances ($\sigma^{2}_{\alpha}=4,\sigma^{2}_{\phi}=110,\rho_{\alpha}=0.740,\rho_{d}=0.026$), for a total of four additional scenarios.
We also extend the same scenario with informative dropout by testing an alternative, neutral dropout mechanism based on a mixed-effects logistic regression model (instead of time to dropout as described in step 2 of Algorithm 1); note that, in this scenario, both models are misspecified. Specifically, we assume that the probability of dropping out during the jthdiscrete period for the kthparticipants in the ithcluster is determined by the following model:
$P_{ijk}(\text{dropout})=\text{logit}^{-1}(\gamma+\nu X_{i}+\omega_{1}\alpha_{i}+\omega_{2}\phi_{ik}),$
where $\gamma=-1.5$ is an intercept term and $\nu,\omega_{1},\omega_{2}$ are parameters equivalent to those of the joint model (that we assume taking the same values described above). The linear predictor (net of the intercept term) is standardized before simulating from this model, and we provide an easy-to-use implementation for simulating data under this dropout mechanism as well at https://github.com/RedDoorAnalytics/simswjm.
@@ 5.5 Performance measures
The key performance measure of interest was bias, which quantifies whether a method targets the true value of a parameter on average. We report bias on both an absolute and relative scale, the latter only for estimands with a true value that is not zero. Additionally, we report empirical and model-based standard errors, to assess the performance of variance estimators and to compare the efficiency of the two methods, and convergence rates and coverage probabilities, for completeness. All performance measures are described in more detail by Morris et al. 45 Finally, note that model fits that did not converge are discarded when computing the performance measures of interest.
@@ 5.6 Number of repetitions
To determine how many repetitions were needed to estimate the performance measures of interest with sufficient precision, we first ran an arbitrary amount of 50 iterations. We then analyzed these preliminary results and concluded that, with 1000 repetitions, we would expect a largest Monte Carlo error for bias of about 2% of the absolute treatment effect value (either of $\delta$ or $\delta_{0},\delta_{1},\delta_{2},\delta_{3}$), which we consider reasonable. This procedure is described in more detail in the supplementary material.
@@ 5.7 Software
The simulation algorithms were implemented in R, while the analysis code for both models (linear mixed and joint model) was implemented in Stata using the built-in gsem command. 50, 42 The simulation study is summarized in R using the rsimsum package. 51 All statistical code to replicate our simulations or fit the joint model using Stata is publicly available on GitHub at https://github.com/ellessenne/swjm, and is archived on Zenodo as well. 52
@@ 5.8 Results
We report the results of the simulation study for the constant intervention model in 5.8.1, and the results for the general time on treatment model in Section 5.8.2.
@@ 5.8.1 Results: simulation study A, constant intervention model
First, we report on the convergence rates of the fitting procedures. Overall, the joint model had lower convergence rates than the mixed model, which was expected given the extra computational complexity; nonetheless, convergence rates were acceptable in all scenarios with non-extreme dropout mechanisms, and we expect such extreme dropout mechanisms (with hazard ratios of 0.5 and 2.0) to be unlikely in practice. These results are described in more detail in Appendix B of the supplementary material available online, and we discuss possible steps to improve convergence rates in practice in Appendix C. According to our experience, we expect these steps to lead to convergence of a joint model in most settings.
Figure 5 depicts the estimated bias for the treatment effect $\delta$ on the longitudinal outcome over repetitions and across all simulation scenarios and models; results are also tabulated in Appendix B of the supplementary material. Overall, the joint model outperformed the linear mixed model, with negligible bias (if any) under every scenario. Conversely, the linear mixed-effects model showed bias estimates in the scenarios with $\nu=-0.2$ and informative dropout, irrespective of the true treatment effect value. These biases were, however, of small magnitude, as relative bias did not exceed 1% (in absolute terms; Appendix B). In all scenarios with $\nu=0.0$ both models yielded unbiased results, and, as expected, both models performed well in all scenarios with non-informative dropout ($\omega_{1}=\omega_{2}=\log(1.0)$).
@@
Interestingly, coverage probability was close to nominal across all scenarios and for both models, possibly because the relative bias to the true treatment effect under the mixed model is not substantial enough to bias statistical inference. These results are reported in Appendix B of the supplementary material.
For the period effects, the linear mixed model showed significant bias at all possible time points when dropout was informative (Appendix B, supplementary material), up to 20% (on a relative scale); the joint model outperformed the linear mixed model, with slight bias (<5%) only in the extreme scenarios with $\omega_{1}=\omega_{2}=\log(0.5),\log(2.0)$. Accordingly, coverage probability was poor when period coefficients were biased for both the linear mixed model and the joint model, and optimal otherwise. As with the treatment effect $\delta$, when dropout was not informative, both models could estimate unbiased period effects with optimal coverage probabilities.
@@
For the ICC, the joint model outperformed the linear mixed model in all scenarios with informative dropout, showing no bias or small bias (at worst, in some extreme scenarios), as in Figure 6; accordingly, coverage was low (Appendix B, supplementary material). The under-coverage is not unexpected because typically, nominal coverage for such second-order parameters may require a large number of clusters, as demonstrated in previous simulations 53. More results on the estimation of the specific variance components are included, once again, in Appendix B of the supplementary material.